OpenStudy (anonymous):

Find maximum and minimum values of y given by equation : x^3+y^3- 3xy =0

OpenStudy (john_es):

You must find the points where the gradient is 0 \[\nabla(x^3+y^3-3xy)=0\] Once you have solved, you will find that this function has a local minimum in \[(x,y)=(1,1)\] with \[f(1,1)=-1\]

OpenStudy (john_es):

You can check that is a true minimum with the Hessian matrix.

OpenStudy (anonymous):

I've found dy/dx = 3x^2 - xy / 3x-3x^2 ..

OpenStudy (anonymous):

therefore y=3x .. no?

OpenStudy (john_es):

I think you should use the partial derivatives, \[\frac{\partial f}{\partial x}=3x^2-3y=0\] \[\frac{\partial f}{\partial y}=3y^2-3x=0\] Where \[f(x,y)=x^3+y^3-3xy\] Solve the system.

OpenStudy (anonymous):

I didn't study that, solve the equation using implicit functions ..

OpenStudy (john_es):

Well, this may be a solution, I hope, \[\frac{d}{dx}(x^3+y^3-3xy)=3x^2+3y^2\frac{dy}{dx}-3y-3x\frac{dy}{dx}=0\] Then extract common factor, dy/dx \[\frac{dy}{dx}=\frac{y-x^2}{y^2-x}\]

OpenStudy (anonymous):

Dy/dx = 0. 0=y-x^2 therefore x=y^1/2 ? then I substitute in the main function?

OpenStudy (john_es):

Exactly, but I would use y=x^2, ans subsitute y for x^2 (easier calculations).

OpenStudy (anonymous):

dude, I stuck :l x^3+x^6-3x^3=0 -2x^3+x^6=0 x^3(-2+x^2)=0 x^3=0 , -2+x^2=0 => x=0, x^2=2 => x=\[\sqrt{2}\]

OpenStudy (john_es):

You have an error in the third line, \[x^3(-2+x^3)=0 \] With this method, \[x=0, x=-\sqrt[3]{2}\]

OpenStudy (john_es):

An error for me too, \[x=0,x=\sqrt[3]{2}\]

OpenStudy (anonymous):

oh , I always do that. thanks for noticing it

OpenStudy (anonymous):

I found y=0.727 .. right?

OpenStudy (john_es):

\[2+y^3-3\sqrt[3]{2}y=0\]It should be the solution from this equation.

OpenStudy (anonymous):

yeah but then?:l root is confusing